HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, all entries  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  134   Thu Apr 3 23:23:20 2025 Zhang Min, Zhou XuAcceleratorBeam injection changeAll the secondary beams are injected into
ESR.
 2025-04-03_23-21-03-095.png 
  133   Thu Apr 3 22:25:49 2025 Usama, KathrinGeneralScreenshots   2025-04-03_22-26-13-116.png2025-04-03_22-26-24-105.png2025-04-03_22-27-26-640.png2025-04-03_22-27-34-598.png2025-04-03_22-28-26-878_tcl1064.tdf 
  132   Thu Apr 3 21:25:40 2025 Usama, KathrinGeneralScreenshots   2025-04-03_21-23-03-733.png2025-04-03_21-23-31-673.png2025-04-03_21-24-01-343_tcl1064.tdf2025-04-03_21-24-35-264.png2025-04-03_21-24-42-340.png 
  131   Thu Apr 3 20:01:58 2025 Usama, KathrinGeneralScreenshots   2025-04-03_20-00-57-730.png2025-04-03_20-02-19-209.png2025-04-03_20-02-55-335.png2025-04-03_20-03-03-078.png2025-04-03_20-31-02-325_tcl1064.tdf 
  130   Thu Apr 3 18:53:02 2025 Usama, KathrinGeneralScreenshots   2025-04-03_18-53-33-868.png2025-04-03_18-54-24-462.png2025-04-03_18-55-11-137.png2025-04-03_18-55-22-991.png 
  129   Thu Apr 3 18:01:49 2025 Ruijiu, CarloAnalysisIdentification of 188Hf   Screenshot_2025-04-03_18-02-52_1.pngsum_1556-1627_spectrum.npzScreenshot_2025-04-03_18-04-19.png 
  128   Thu Apr 3 17:49:57 2025 Usama, KathrinGeneralScreenshots   2025-04-03_17-49-01-753.png2025-04-03_17-50-33-265.png2025-04-03_17-51-26-693.png2025-04-03_17-52-07-151.png 
  127   Thu Apr 3 17:38:47 2025 Shahab, Zac, Carlo, Esther and many moreDAQRSAs run again!first file name RSA01....17.38 onwards.  IMG_4972.JPGIMG_4973.JPG 
  126   Thu Apr 3 17:01:07 2025 CJ Griffin, M Cockshutt, G LeckenbyGeneralMoving PLEIADES detectors inThe inner PLEIADES is moved into position
close to the beam.
Device control for the outside pocket is
  
  125   Thu Apr 3 16:44:18 2025 Usama, KathrinGeneralScreenshots   2025-04-03_16-44-51-040.png2025-04-03_16-45-58-954.png2025-04-03_16-46-46-981.png2025-04-03_16-48-47-815.png 
  124   Thu Apr 3 15:57:16 2025 Shahab, Yue, Zac, EstherDAQData acquisitions runningAll data acquisitions (both RSAs and NTCAP)
are now running.
  
  123   Thu Apr 3 15:55:04 2025 Shahab, Yue, EstherCalibrationSchottky settingsSettings similar to the 2-photon experiment.

Central frequency: 411.742243 MHz (based
  
  122   Thu Apr 3 15:21:42 2025 CJGGeneralPLEIADES shift instructionsPLEIADES shift instructions available at:
https://docs.google.com/document/d/19b4wyugC3OpVAjwt1ONM9u8EYdMJsBpL7GVNzoziv0M/edit?usp=sharing
  
  121   Thu Apr 3 13:36:17 2025 Ruijiu,Yuri, Wolfram,ZsoltAnalysis188Hf identification   Screenshot_2025-04-03_15-19-47.pngScreenshot_2025-04-03_15-21-53_1.pngScreenshot_2025-04-03_15-23-37.pngsimulation_result.outScreenshot_2025-04-03_15-41-11.png 
  120   Thu Apr 3 13:32:42 2025 CN, IDGeneralScreenshots   2025-04-03_13-31-16-104.png2025-04-03_13-31-21-711.png2025-04-03_13-31-42-214.png 
  119   Thu Apr 3 11:51:00 2025 CN, IDDAQRSA started for 188Hf ESR tuningStarted both RSAs for the tuning of the ESR
to 188Hf. Both pockets are in parking position
OUTSIDE.
  
  118   Thu Apr 3 10:07:34 2025 CN, ZP, CJG, MCGeneralScreenshots   2025-04-03_10-04-14-352.png2025-04-03_10-04-23-381.png2025-04-03_10-05-06-655.png 
  117   Thu Apr 3 09:53:02 2025 Ruijiu Chen,WolframAnalysisGood injection ratioCurrently, we only use 37% of the injections
in our data analysis. However, 63% of the
injections have a height below 100, and these
 photo_2025-04-03_09-53-26.jpg 
  116   Thu Apr 3 09:43:14 2025 AllAcceleratorStop 194Pb, switch to 208Pb82+The repair of SIS electron cooler power supplies

In the meantime we received a few injections
  
  115   Thu Apr 3 09:34:42 2025 IDGeneralFinished in the ESR as of 9:05We had stopped the RSAs at 8:00 (call for
restart)
Time stamp computer stopped and started at
  
  114   Thu Apr 3 08:02:04 2025 Team PLEIADES (JG, ID, CJG, MC)DetectorsEntering ESR for PLEIADESWe are going to enter the ESR to check on
the PLEIADES pocket detectors and install
a nitrogen cooling.
  
  113   Thu Apr 3 07:59:09 2025 ZP, CN, IDGeneralScreenshotsScreenshots before opening ESR.  2025-04-03_07-58-00-874.png2025-04-03_07-57-08-478.png2025-04-03_07-57-36-781.png 
  112   Thu Apr 3 06:56:31 2025 JEL, ZNRunsScreenshots   2025-04-03_06-55-08-030.png2025-04-03_06-54-58-081.png2025-04-03_06-55-25-792.png 
  111   Thu Apr 3 06:33:08 2025 Zac, JohanGeneralBeam lost at 06:27, returned at 06:34Spike in beam intensity at 06:27 (5x usual
intensity) and then beam is lost. HKR says
it's a problem with the magnet's power supply.
  
  110   Thu Apr 3 06:16:32 2025 JEL, ZNGeneralScreenshotsStable since last entry  2025-04-03_06-15-12-250.png2025-04-03_06-15-00-458.png2025-04-03_06-14-54-153.png 
  109   Thu Apr 3 05:11:37 2025 Zac, JohanGeneralBeam lost at 05:07, returned at 05:14    
  108   Thu Apr 3 03:48:40 2025 JEL, ZNRunsScreenshots
The peaks in the analyzer are slowly shifting
back to the original position. 
 2025-04-03_03-47-28-530.png2025-04-03_03-47-11-337.png2025-04-03_03-47-42-665.png 
  107   Thu Apr 3 02:59:15 2025 JEL, ZNRunsScreenshot, beam received, baem lost, beam back againBeam is back for 10 mins-

The peaks on the analyzer appear shifted
 2025-04-03_02-58-06-109.png2025-04-03_02-57-54-019.png2025-04-03_02-57-31-430.png 
  106   Thu Apr 3 02:52:27 2025 Zac, JohanGeneralBeam is back at 02:52    
  105   Thu Apr 3 02:01:29 2025 Zac, JohanGeneralQuadrupole issuesHKR has called someone to fix the quadrupole,
they think it will take about an hour.
  
  104   Thu Apr 3 01:56:57 2025 Zac, JohanGeneralProblems with one of the quadrupolesSince around 01:45 we get spills but see nothing
on analysers. HKR says there is a problem
with one of the quadrupoles; they are working
  
  103   Thu Apr 3 01:29:49 2025 Zac, JohanGeneralBeam is back at 01:27    
  102   Thu Apr 3 01:19:44 2025 Zac, JohanGeneralBeam fluctuationsBeam gone again between 01:14 and 01:18, then
again at 01:19. HKR says there are problems
with the source.
  
  101   Thu Apr 3 01:14:16 2025 Zac, JohanGeneralBeam is back at 01:12    
  100   Thu Apr 3 00:49:00 2025 Zac, JohanGeneralBeam lost at 00:46    
  99   Wed Apr 2 23:40:23 2025 JEL, ZNGeneralScreenshots taken just as beam came back
We are seeing spills being received but no
peaks are seen in the analyzers. 
 2025-04-02_23-39-29-655.png2025-04-02_23-37-10-651.png2025-04-02_23-36-58-664.png 
  98   Wed Apr 2 23:33:43 2025 ZacAcceleratorBeam is gone at 23:29    
  97   Wed Apr 2 23:27:35 2025 Zac, JohanAcceleratorBeam is back 23:25    
  96   Wed Apr 2 23:15:52 2025 Zac, JohanAcceleratorNo injections from SISNight shift starts with no injections from
SIS! We still have triggers; we called HKR
and are waiting for response.
  
  95   Wed Apr 2 22:47:25 2025 ID, HA (Radio Shahab as background)GeneralLassie and Schottky TV screentshotsWe managed to take some screenshots while
Shahab harrassed us with AI-created "RammstAIn"
music... happy to end my shift now!
 2025-04-02_22-44-18-420.png2025-04-02_22-45-12-849.png2025-04-02_22-55-14-900.png 
  94   Wed Apr 2 21:22:41 2025 ID, HAGeneralLassie Come-Home Screenshots (wuff!)(Who let the dog out?)
Sharing with FRS 6:1 since a while now.
 2025-04-02_21-25-00-716.png2025-04-02_21-25-42-535.png2025-04-02_21-26-12-022.png 
  93   Wed Apr 2 20:57:34 2025 ShahabDetectors10GB cables for the DAQ rackWe have been working together with IT on the
speed of data transfer: with the new 10 GB
interface we can now transfer data from NAS
  
  92   Wed Apr 2 20:15:29 2025 ID, HAGeneralLassie screenshots + Schottky TVScreeeeeeeenshooooots!  2025-04-02_20-15-48-827_tcl1064.png2025-04-02_20-17-14-017.png2025-04-02_20-18-29-150.png 
  91   Wed Apr 2 19:13:50 2025 ID, HAGeneralLassie screenshots and Schottky screenshotLooks a bit more regular now (FRS people having
dinner???)
 2025-04-02_19-10-29-388_tcl1064.png2025-04-02_19-14-53-428_tcl1146.png2025-04-02_19-19-23-765.png 
  90   Wed Apr 2 18:31:52 2025 IDGeneralFYI for morning shifts ThuThe PLEIADES team (pocket detectors) has discussed
with Wolfram to get access to the ESR tomorrow
(Thu) morning ~8:00 to check their detectors
  
  89   Wed Apr 2 18:22:30 2025 IDGeneralSpill cycle distributionThe spill cycle distributions is absolutely
nuts since beginning of the shift. On the
positive side: we got a lot of spills since
  
  88   Wed Apr 2 18:15:06 2025 ID, HA, MWGeneralLassie screenshots   2025-04-02_18-14-21-624.png 
  87   Wed Apr 2 18:13:20 2025 ID, HAGeneralRSA spectra   IMG_0087.jpeg 
  86   Wed Apr 2 17:46:12 2025 Ruijiu, Wolfram, Xinliang, WangqianAnalysisThe spectrum of 194Pb82+ vs 184Pt78+In summary, the statistical results collected
between 5:00 and 12:00 AM on April 2, 2025.
For each injection, we calculate the frequency
 photo_2025-04-02_17-47-24.jpgplot_fft.cxxScreenshot_2025-04-02_22-00-57_1.png 
  85   Wed Apr 2 17:06:45 2025 ID, HA et alGeneralLassie screenshots   2025-04-02_17-09-17-201.png 
  84   Wed Apr 2 16:28:14 2025 SS, ID, HA etcDAQNTCAP stop/startNTCAP was stopped and restarted.   
  83   Wed Apr 2 15:46:24 2025 ID, HA,MWGeneralBeam back in ESR, beam is back, we started DAQ again. All files
created after 15:45 should be good from now
on.
 20250402_154135.jpg 
  82   Wed Apr 2 15:07:52 2025 Michael, Alex, MaeveAcceleratorStill not a good beam in the ESRThe (power supply of) SIS kicker magnet has
been repaired and the beam has been restored
from about 14:20.
  
  81   Wed Apr 2 13:35:17 2025 Michael, Alex, MaeveDAQRSAs data takingSince there will be no beam for about one
hour, the data taking in RSAs has been stopped.
  
  80   Wed Apr 2 13:27:52 2025 Michael, Alex, MaeveAcceleratorNo beam in the ESRThere is no beam in the ESR since couple of
minutes. According to the info from HKR,
the experts started to repair the magnets
  
  79   Wed Apr 2 12:58:15 2025 Ruijiu, WolframAnalysisfrequency drift   photo_2025-04-02_15-24-34.jpgphoto_2025-04-02_15-28-44.jpgfrequency_194Pb82.xlsx 
  78   Wed Apr 2 12:13:57 2025 Michael, AlexAcceleratorSome empty injections triggered while beam patterns are changedOperators seem to play with spill cycles,
we get every 1:1 shot, then 1:10 again. During
this, we noticed a few "empty injections"
  
  77   Wed Apr 2 11:09:42 2025 Michael, AlexAcceleratorThe beam is back in the ESRWe have the beam again the ESR since around
11:00.
The Schottky spectra look the same as before.
  
  76   Wed Apr 2 10:55:16 2025 Michael, Alex, WolframAcceleratorNo beam in the ESRSince few minutes do not see any beam in the
ESR (on RSAs). There is beam in SIS and injection
requests from the ESR seem to be there as
  
  75   Wed Apr 2 10:31:15 2025 Michael, AlexAcceleratorThe beam is back from 10:10     
  74   Wed Apr 2 10:05:06 2025 Michael, Alex, YuriAcceleratorNo beam since 09:59The SIS kicker has failed again.   
  73   Wed Apr 2 10:02:40 2025 Sergey + YuriAcceleratorESR TimeoutAt around 9:40, ESR timeout changed from 10
s to 30 s
Triggering only on real injections.
  
  72   Wed Apr 2 09:28:59 2025 Michael, AlexGeneralEmpty injections in the ESRSince the FRS started to take the beam again,
the ESR gets 1/10 shots from SIS. But, the
ESR seems to request about twice more frequently
  
  71   Wed Apr 2 08:47:27 2025 Michael, AlexGeneralThe beam is backThe beam is back. But, the experts have found
a problem with one of the modules of the
SIS kicker magnet. And, currently they have
  
  70   Wed Apr 2 08:04:47 2025 Michael, AlexGeneralNo beam The SIS kicker has failed again at 08:01.
Since this has happened already few times,
the experts from the on-call duty service
  
  69   Wed Apr 2 07:41:10 2025 Michael, AlexGeneralFrequency shift in SchottkyAfter the beam  was restored at 07:25, we
think that the frequency peak in the 244
MHz RSA2 has jumped again (closer to the
  
  68   Wed Apr 2 07:39:41 2025 Michael, AlexGeneralThe beam is back at 07:35    
  67   Wed Apr 2 07:33:59 2025 Michael, AlexGeneralBeam gone again7:29, beam gone again, operators say they
need another 5 minutes.
  
  66   Wed Apr 2 07:25:15 2025 Carlo, Chris, Alex, Michael GeneralBeam back07:20  we have beam back.   
  65   Wed Apr 2 07:16:33 2025 Carlo, Chris, Alex, MichaelDAQESR not requesting7:05 ESR not requesting new pulses from the
SIS. 
Operators are trying to fix the problem.
  
  64   Wed Apr 2 05:48:31 2025 Chris, Carlo, ZacGeneralBeam returnedBeam was returned to the ESR at 05:11.

HKR said there had been some problem with
  
  63   Wed Apr 2 04:59:50 2025 Zac, Chris, CarloGeneralReceiving injections but not seeing anything on RSAsStarting at around 04:09 we don't see anything
on RSAs. RSAs are still triggering and writing
files, but only see noise. Lassie indicates
 2025-04-02_05-18-32-212_tcl1146.png2025-04-02_05-21-50-973_tcl1146.png 
  62   Wed Apr 2 04:33:57 2025 Carlo, Zac, ChrisDAQScreenshots Lassiescreenshots of Lassie  2025-04-02_04-28-59-944_tcl1064.png2025-04-02_03-38-54-877_tcl1064.png 
  61   Wed Apr 2 02:51:31 2025 Carlo, Zac, ChrisDAQBeam is back02:44 beam is back   
  60   Wed Apr 2 01:46:52 2025 Chris, Carlo, ZacGeneralBeam lossHKR just called to say a power supply to the
SIS18 extraction kicker has failed.
The service person has been called and is
  
  59   Wed Apr 2 00:59:43 2025 Zac, Carlo, ChrisAcceleratorLost the beamBeam was lost from 00:44 until 00:59   
  58   Wed Apr 2 00:08:57 2025 Ruijiu Chen,ShahabAnalysisthe peak corresponding to 194Pb⁸²⁺ was mixed with signals from some unknown ions.I examined the spectrum of 194Pb at the 211
harmonic. It was observed that the peak corresponding
to 194Pb82+ was mixed with signals 
 photo_2025-04-02_00-12-38.jpgphoto_2025-04-02_00-12-43.jpg 
  57   Tue Apr 1 23:56:18 2025 Shahab, Carlo, Zac, Yue, MichaelDAQChange of RSA01 settings for new harmonic=212.23:55 stop the RSAs. 

410:
  
  56   Tue Apr 1 23:23:18 2025 C GriffinGeneralPLEIADES removedLeakage currents were steadily increasing
in the PLEIADES detectors so we decided to
stop taking data, move both the detectors
  
  55   Tue Apr 1 23:07:15 2025 ShahabDAQshort interruption of queqiao script for RSAshas been restored.   
  54   Tue Apr 1 22:53:59 2025 EstherGeneralLassie exports   2025-04-01_22-53-17-548_tcl1146.png2025-04-01_22-53-21-680_tcl1146.tdf 
  53   Tue Apr 1 22:05:24 2025 Ruijiu Analysisanalysis resultThis is a preliminary result during 4:00-19:00
April 1st 2025. 
I correct the frequency drift of 194Pb82+
 202504012157.png202504012157.pdf202504012206.pngphoto_2025-04-01_22-43-41.jpg 
  52   Tue Apr 1 22:03:07 2025 Jeroen, Zac, Esther, Shahab, WolframDAQOptimisation of settingsEstimation of potential gain by changing.
Files from this points are going to rubbish.
Potential gains are c. 20% 
  
  51   Tue Apr 1 21:05:20 2025 Jeroen, Zac, Esther, Shahab, WolframDAQChange of settings for shorter acq. window> Acquisition stopped at 20:36. 

> First file from this setting: RSA01-2025.04.01.03.52.59.247.tiq
  
  50   Tue Apr 1 20:37:04 2025 JeroenDAQChange of settings for shorter acq. windowAcquisition stopped at 20:36. 

First file from this setting: RSA01-2025.04.01.03.52.59.247.tiq
  
  49   Tue Apr 1 20:19:23 2025 Jeroen, Zac, EstherAcceleratorBeam fell out> > No beam from 19:57:41.

> Beam back from 20:02
 2025-04-01_20-17-10-651_tcl1064.png 
  48   Tue Apr 1 20:02:19 2025 Jeroen, Zac, EstherAcceleratorBeam fell out> No beam from 19:57:41.

Beam back from 20:02
  
  47   Tue Apr 1 20:00:14 2025 Jeroen, Zac, EstherAcceleratorBeam fell outNo beam from 19:57:41.   
  46   Tue Apr 1 17:08:13 2025 JG, ID, MC, CGDetectorspleiades in positionThe pleiades detectors are now moved into
a position which does not disturb the stored
beam. 
  
  45   Tue Apr 1 16:47:32 2025 Ruijiu, Jan, ShahabAnalysisThe frequency of 194Pb82+ is changed to 408.600 MHz   Screenshot_2025-04-01_16-46-56.pngphoto_2025-04-01_16-54-20.jpg 
  44   Tue Apr 1 16:44:32 2025 Jeroen, EstherGeneralNo SIS beam since 14:35Beam resumed at 16:40.   2025-04-01_16-47-13-677_tcl1146.png2025-04-01_16-47-10-870_tcl1146.tdf 
  43   Tue Apr 1 15:38:50 2025 ruijiu chenAnalysisfrequency drift   Screenshot_2025-04-01_15-38-26_1.pngScreenshot_2025-04-01_15-41-47.png 
  42   Tue Apr 1 15:32:59 2025 ZacDAQNTCAP RestartedNTCAP Has been restarted (we restart every
day around 15:00 so data goes in a new folder)
  
  41   Tue Apr 1 14:54:24 2025 allGeneralNo SIS beam since 14:35    
  40   Tue Apr 1 14:25:21 2025 ruijiu AnalysisThe frequency of 194Pb82+ has been changed from 408.606MHz to 408.602MHy during last 12 hoursThe following images display the frequency
spectrum at the start of every hour from
3:00 to 14:00.
 12x 
  39   Tue Apr 1 14:23:01 2025 JGGeneralPleiades pocket position testsWe started playing with the detector pockets
of the pleiades arrays at 14:23 01.04.2025.
  
  38   Tue Apr 1 14:17:07 2025 RuijiuAnalysisPID of 194Pb82+ 2025.04.01.14.13.07datafile = "/data.local/G22-00018_00203/results/analyzers/RSA01/RSA01-2025.04.01.14.13.07.068.tiq_spectrum.npz"
filep = "/data.local/G22-00018_00203/scripts/OnlineDataAnalysisSystem/rionid_exercises/exercise_194Pb_Exp_2025/New_TEline2022-ESR-194Pb-208Pb-nostripper_524MeV.lpp"
alphap = "1.399"
 Screenshot_2025-04-01_14-16-42.png 
  37   Tue Apr 1 14:02:23 2025 Jelena B, SiddharthGeneralLassie   2025-04-01_14-01-37-464_tcl1064.png 
  36   Tue Apr 1 12:58:27 2025 Jelena B, SiddharthGeneralLassie   2025-04-01_12-57-50-275_tcl1064.png 
  35   Tue Apr 1 11:47:14 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_11-46-35-517_tcl1064.png    2025-04-01_11-46-35-517_tcl1064.png 
  34   Tue Apr 1 10:34:20 2025 Jelena B, SiddharthGeneralResolved issue with SIS18Beam is back with some fluctuations.
We did not have beam from 9:34 until 10:08,
around 35 minutes. 
 2025-04-01_10-33-38-948_tcl1064.png 
  33   Tue Apr 1 10:03:56 2025 Jelena B, SiddharthGeneralTechnical issue with SIS18 From HKR info: there is a technical issue
with SIS18 and we will have no beam for next
25 minutes. 
 2025-04-01_10-02-16-454_tcl1064.png 
  32   Tue Apr 1 08:36:25 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_08-37-32-155_tcl1064.png   2025-04-01_08-37-32-155_tcl1064.png 
  31   Tue Apr 1 06:53:51 2025 Jelena B,SiddharthGeneralLassie Monitor   2025-04-01_06-49-59-989_tcl1064.png 
  30   Tue Apr 1 06:31:46 2025 MichaelGeneralLassie Monitor 2025-04-01_06-31-10-211_tcl1064.png   2025-04-01_06-31-10-211_tcl1064.png 
  29   Tue Apr 1 05:58:51 2025 MichaelGeneralLassie Monitor 2025-04-01_05-58-09-122_tcl1064.png   2025-04-01_05-58-09-122_tcl1064.png 
  28   Tue Apr 1 05:29:45 2025 MichaelGeneralLassie Monitor 2025-04-01_05-29-01-845_tcl1064.png   2025-04-01_05-29-01-845_tcl1064.png 
  27   Tue Apr 1 04:59:14 2025 MichaelGeneralLassie Monitor 2025-04-01_04-58-31-351_tcl1064.png   2025-04-01_04-58-31-351_tcl1064.png 
  26   Tue Apr 1 04:27:10 2025 MichaelGeneralLassie Monitor 2025-04-01_04-25-53-171_tcl1064.png   2025-04-01_04-25-53-171_tcl1064.png 
  25   Tue Apr 1 03:55:59 2025 MichaelGeneralLassie Monitor 2025-04-01_03-55-21-183_tcl1064.png   2025-04-01_03-55-21-183_tcl1064.png 
  24   Tue Apr 1 03:52:56 2025 Shahab, Michael, Yue, Zac, CarloDAQgood filesall settings are done.
any files coming after this one is good:
  
  23   Tue Apr 1 03:50:12 2025 Shahab, Yue, Carlo, Zac, Esther, MichaelDAQTuners of the both cavities set410 MHz cavity resonance frequency tuned.
motors both on 50 mm
245 MHz cavity resonance frequency tuned.
  
  22   Tue Apr 1 03:27:44 2025 MichaelGeneralLassie Monitor 2025-04-01_03-22-05-395_tcl1064.png    2025-04-01_03-22-05-395_tcl1064.png 
  21   Tue Apr 1 02:55:06 2025 Shahab, Michael, Yue, Zac, CarloDAQnew RSA recordinganything from RSA01-2025.04.01.02.54.25.976.tiq
onwards is good data
 IMG_4955.JPGIMG_4956.JPGIMG_4954.JPG 
  20   Tue Apr 1 02:37:28 2025 H. Weick, S. LitvinovAcceleratorIsoc-setting +scrapersFirst a setting with orbit bump for injection
on central orbit was tried.
- quadrupoles tuned and scrapers optimized
 IMG_20250401_013626.jpgIMG_20250401_013636.jpgIMG_20250401_013646.jpgIMG_20250401_023358.jpg 
  19   Tue Apr 1 02:20:36 2025 ruijiuAnalysischange back the frequency of 194Pb82frequency of 194Pb82+ at 211 408606000
frequency of 194Pb82+ at 126 244001000
 Screenshot_2025-04-01_02-20-09.png 
  18   Tue Apr 1 02:03:33 2025 ruijiuAnalysisThe frequency of 194Pb82+ has been changed from 408.522 MHz to 408.637 MHzthis is the new setting after we lost the
old one.
 Screenshot_2025-04-01_02-03-14.png 
  17   Tue Apr 1 02:01:06 2025 Zac, Shahab, Ruijiu, Yuri, Sergey, Wolfram, Carlo, MichaelAcceleratorSettings lostSettings lost after the beam was lost. Trying
to recover particle identification now that
beam has returned.
  
  16   Mon Mar 31 23:34:58 2025 ruijiu, shahab, yuri, wolfram, sergeyAnalysisidentificaton of 194Pb82+datafile = "/data.local/G22-00018_00203/results/analyzers/RSA01/RSA01-2025.03.31.22.18.16.965.tiq_spectrum.npz"
filep = "/data.local/G22-00018_00203/scripts/OnlineDataAnalysisSystem/rionid_exercises/exercise_194Pb_Exp_2025/New_TEline2022-ESR-194Pb-208Pb-nostripper_524MeV.lpp"
alphap = "1.4"
 8x 
  15   Mon Mar 31 22:35:36 2025 ruijiu ChenAnalysisIdentification of 194Pb82   7x 
  14   Mon Mar 31 22:13:04 2025 Shahab, Yue, Carlo, Zac, EstherRunsPID continuedcontinuing PID with better / higher resolution:

RSA01-2025.03.31.21.10.54.758.tiq_spectrum.npz
 photo_2025-03-31_22-17-30.jpg 
  13   Mon Mar 31 22:03:56 2025 Shahab, Yue, Carlo, Zac, EstherRunsPIDtrying to do PID with some injections:

RSAs are still on winter time. they show
  
  12   Mon Mar 31 20:53:02 2025 Shahab, Carlo, Zac, Esther, MichaelDAQDAQ start20:28 started NTCAP
started queqiao file transfer for RSAs only
started analyzer looper for RSAs
  
  11   Mon Mar 31 20:42:38 2025 H. Weick, S. LitvinovAcceleratorfragment setting, E-SIS~ 20:00 changed E-SIS to 524.0 MeV/u to get
194Pb at the Brho of the primary beam before
(=7.2008 Tm).
  
  10   Mon Mar 31 17:46:21 2025 EstherDAQNTCAP scaler assignmentThe setup contains two NI-6612 I/O cards with
BNC-2121 connector accessories that provide
eight inputs each for TTL scaler signals.
  
  9   Mon Mar 31 16:53:03 2025 ShahabDAQstop and delete NTCAP data from Weekend teststop and delete NTCAP data from Weekend test

for now the NAS data2 bank has been flushed
  
  8   Mon Mar 31 16:05:48 2025 ruijiu, shahab, yuyue, zac, carlo, erstherAnalysischeck the trigger signal in the scaler data of NTCAP1. convert copy a tdms file into the folder:
(pyroot) gpuuser@appc256:/data.local/G22-00018_00203/data/ntcap/sc/SC_2025-03-28_18-48-51$
cp /mnt/ntcap-data2/G22-00018_00203/data/ntcap/sc/SC_2025-03-28_18-
 Screenshot_20250331_161004.pngScreenshot_2025-03-31_16-23-55.png 
  7   Mon Mar 31 14:59:18 2025 ruijiuAnalysisrun rionidgui program at HKRAt HKR, there is a Windows computer equipped
with X2go software, which allows users to
log in to the GPU computer. The login process
 IMG_20250331_150222.jpgIMG_20250331_150216.jpgIMG_20250331_151940.jpgIMG_20250331_151929.jpgIMG_20250331_151959.jpg 
  6   Mon Mar 31 10:46:43 2025 Ruijiu ChenAnalysisThis program is used to combine spectrum which is recored by RSA into one spectrum.This program is used to combine spectrum of
many injections into one spectrum.
 generate_data_1.sh 
  5   Fri Mar 28 22:20:47 2025 Sergey, ShahabGeneralPatternthe ACC pattern for both experiments.  pattern.JPG 
  4   Fri Mar 28 19:42:20 2025 Ruijiu ChenAnalysisQueqiao Program for RSA DataThe `queqiao` program used for RSA data analysis
is located at:
  
  
  3   Thu Mar 27 17:20:40 2025 Esther, Yue, Carlo, Zac, ShahabDetectorsSetup GenesysWe set up the Genesys. Channel map:

tr0:
 2025-03-27_17-19-33-724.png 
  2   Thu Mar 27 09:42:04 2025 Ruijiu ChenSimulationsSimulation Results for 194Pb and 188HfSummary:

1.Simulated spectra for 194Pb and 188Hf.
 9x 
  1   Mon Mar 24 14:42:33 2025 Jan GloriusCollaborationProposals Attached are the proposals submitted to and
approved by G-PAC.
 G-22-00203-UR.pdfG-22-00018-1.pdf 
ELOG V3.1.5-fc6679b